# CAML: Collaborative Auxiliary Modality Learning for Multi-Agent Systems

### v1

@@ A.6 Limitations and Future Work
Even though the advancements of CAML, there are some limitations. One limitation is that if the modalities are misaligned, the model may struggle to perform effective fusion, leading to incorrect predictions. The auxiliary modalities or views from collaborative agents may become noise, useless or even degrading performance. Another limitation is the increasing system complexity. As the number of agents increases, the complexity of the system grows. The fusion of multi-agent and multi-modal data introduces challenges related to coordination overhead, which may lead to delays in the collaborative learning process. Future work can focus on address these limitations.

### v2

@@ A.6 Limitations and Future Work
CAML can handle different and mixed modalities for the agents at test time. Feature embeddings are first extracted by modality-specific encoders and then fused across agents (e.g., via concatenation or cross-attention) for downstream decision-making. In the student model of CAML, the modalities for each agent can be configured individually. The agents do not all need to have the same modalities. The current student model training does not dynamically change the modalities over time. However, we can handle dynamic changes without changing the architecture. Our planned solution is to incorporate a modality dropout strategy. During training of knowledge distillation, we randomly drop out or mask some modality for each agent in the student model. In this way, the student model learns to handle cases where some modalities may vary dynamically for different agents. Exploring this strategy would be a promising direction for future work. Also similar to prior works (e.g., STGN [8], Who2com [22], When2com [21], and Where2comm [17]), CAML assumes reliable communication during training. While this is common in the field, future work could explore robustness under stochastic communication delay or failure.